OpenStudy (anonymous):

If f(x) =3x+4,find f^-1(x)

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

na

OpenStudy (anonymous):

Good luck then I guess.

OpenStudy (epoweritheta):

batman u r cool

OpenStudy (anonymous):

Thanks lol

OpenStudy (anonymous):

\[f(x) = 3x+4\] To find the inverse, follow the steps I've posted above. So, first thing replace f(x) with y. \[y=3x+4\] Our second step is to switch the x's and y's around, so: \[x=3y+4\] Now we just solve for y \[x-4=3y \] \[y = \frac{ x-4 }{ 3 }\] and finally we replace y with f^-1(x) \[f ^{-1}(x) = \frac{ x-4 }{ 3 }\] Do you need the reason why we do this or just this?

OpenStudy (anonymous):

\[f ^{-1}(x) \neq \frac{ 1 }{ f(x) }\] the negative 1 does not represent an exponent here.